OpenStudy (anonymous):

A water ride at a local water park has a ride shaped like a cone that acts like a funnel whereby guests swirl around the cone until they drop through its center. There is one ride for adults and a similar, smaller version for children. If the adult ride has a radius of 25 feet and the child ride has a radius of 10 feet, what is the ratio between the volumes of each ride?

OpenStudy (mertsj):

Oh no. Not this question again. What was wrong with the previous answer provided?

OpenStudy (mertsj):

Are we to assume that the heights are the same?

OpenStudy (anonymous):

The last question like this had different numbers.

OpenStudy (anonymous):

I really have no idea....I just know you are supposed to put them in ratios

OpenStudy (anonymous):

25:4 15:6 125:8 5:2 these are the answer choices

OpenStudy (mertsj):

Since we are given no information to the contrary, let us assume that the height is constant.

OpenStudy (mertsj):

\[\frac{V _{1}}{V _{2}}=\frac{\frac{1}{3}\pi (25)^2h}{\frac{1}{3} \pi (10^2)h}=\frac{625}{100}=\frac{25}{4}\]

OpenStudy (anonymous):

Okay, I thought that was the answer but I wasn't sure. Thanks :)

OpenStudy (mertsj):

yw